OpenStudy (darkigloo):

lim when x approaches a (√x - √a) / x-a I did this: (√x - √a) / x-a = (x+a) / ( (x-a)(√x - √a) ) = (a+a) / ( (a-a)(√a - √a) ) = 2a/0 = Does not exist But I'm not sure I did that right.

myininaya (myininaya):

instead factor the denominator.

myininaya (myininaya):

\[x-a=(\sqrt{x}-\sqrt{a})(\sqrt{x}+\sqrt{a})\]

myininaya (myininaya):

I just treated it like it was a difference of squares basically

myininaya (myininaya):

if you don't like that way. Relation the numerator. You sorta did it incorrectly above.

OpenStudy (darkigloo):

i did the way you said an i got 2√a

myininaya (myininaya):

one over that right?

OpenStudy (darkigloo):

oh yeah.